Patient from Colombia comes into the CVS Pharmacy and has a reading on the meter : 3.5 millimolar .1 mMole Glucose = 180 mg Glucose. The conversion to the American measurements is :

a) 63 mg/100 mL

b) 196 mg/dL

c) 261 mg / 100 mL

d) 2610 mg / 100 mL

e) 630 mg /100 mL
Answer #1

Molarity = mol/L.

1 milimolar solution means 1 milimol solute in present 1 L solution.

Then , 3.5 milimolar glucose = 3.5 milimol glucose in 1 L solution.

Now, 3.5 milimol glucose = 3.5*180= 630 mg glucose.

[ Molar mass of glucose = 180 g/mol]

So, 630 mg glucose is dissolved in 1000 mL solution.

Or, 63 mg glucose is dissolved in 100 mL solution.

Therefore, 3.5 milimolar glucose solution = 63mg/100mL

So, correct option is a.